OpenStudy (anonymous):

Use mathematical induction to prove the statement is true for all positive integers n, or show why it is false. @SithsAndGiggles 1^2 + 4^2 + 7^2 + ... + (3n - 2)^2 = n(6n^2-3n-1)/2

OpenStudy (anonymous):

Okay I don't see a readily available counter-example. According to WA, the formula is correct, but you want to prove it's correct. An induction proof consists of the following: (1) base case (2) induction step The base case establishes that the formula, in this case \[\sum_{i=1}^n(3i-2)^2=\frac{n(6n^2-3n-1)}{2}\] holds for the first (base) index. What this means is we plug in \(n=1\): \[\sum_{i=1}^1(3i-2)^2=(3\cdot1-2)^2=1^1=1\\ \frac{1(6\cdot1^2-3\cdot1-1)}{2}=\frac{6-3-1}{2}=\frac{2}{2}=1\] So the base case holds. So far so good?

OpenStudy (anonymous):

yes! im folowing

OpenStudy (anonymous):

so thats it right?

OpenStudy (anonymous):

oh noo theres the induction step

OpenStudy (anonymous):

The induction step involves this reasoning: If the formula holds for an arbitrary \(n\), and if we can prove it holds for the next \(n\), then the formula holds in general. In other words, if we assume the formula is correct for \(n=k\), and we use that assumption to show it holds for \(n=k+1\), then the formula is true for all \(n\).

OpenStudy (anonymous):

In context: Assume the formula holds for \(n=k\). This means \[1^2+4^2+\cdots+(3k-2)^2=\frac{k(6k^2-3k-1)}{2}\] What we want to show: \[1^2+4^2+\cdots+(3k-2)^2+(3(k+1)-2)^2=\frac{(k+1)(6(k+1)^2-3(k+1)-1)}{2}\] Notice that we just add the \(k+1\)-th term to the left side and substitute \(k+1\) for \(n\) on the right side. Do you follow?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the ending -3(k+1-2? it cut off

OpenStudy (anonymous):

so that's it? it is true? @SithsAndGiggles

OpenStudy (anonymous):

No, we haven't shown anything yet. I'm only making the connection between the induction proof strategy and the actual facts we have to work with.

OpenStudy (anonymous):

ohhh now i understand, i got a little confused

OpenStudy (anonymous):

As for what was cut off: \[1^2+4^2+\cdots+(3k-2)^2+(3(k+1)-2)^2\]is equal to\[\frac{(k+1)(6(k+1)^2-3(k+1)-1)}{2}\]

OpenStudy (anonymous):

To be precise, we want to prove it's equal.

OpenStudy (anonymous):

To do that, we start with the left hand side: \[\color{red}{1^2+4^2+\cdots+(3k-2)^2}+(3(k+1)-2)^2\] Recall the assumption for the first \(k\) terms: \[1^2+4^2+\cdots+(3k-2)^2=\frac{k(6k^2-3k-1)}{2}\] Substituting, we have \[\color{red}{\frac{k(6k^2-3k-1)}{2}}+(3(k+1)-2)^2\]

OpenStudy (anonymous):

We want to be able to combine the terms so we can establish \[\color{red}{\frac{k(6k^2-3k-1)}{2}}+(3(k+1)-2)^2=\frac{(k+1)(6(k+1)^2-3(k+1)-1)}{2}\]

OpenStudy (anonymous):

It's all algebra from here on out. \[\begin{align*}\frac{k(6k^2-3k-1)}{2}+(3(k+1)-2)^2&=\frac{1}{2}\left(k(6k^2-3k-1)+2(3(k+1)-2)^2\right)\\ &=\frac{1}{2}\left(6k^3-3k^2-k+2(3k+1)^2\right)\\ &=\frac{1}{2}\left(6k^3-3k^2-k+2(9k^2+6k+1)\right)\\ &=\frac{1}{2}\left(6k^3-3k^2-k+18k^2+12k+2)\right)\\ &=\frac{1}{2}\left(6k^3+15k^2+11k+2\right) \end{align*}\] So to sum up, we have to prove this equality: \[\frac{1}{2}\left(6k^3+15k^2+11k+2\right)=\frac{(k+1)(6(k+1)^2-3(k+1)-1)}{2}\] Notice we already have the factor of \(\dfrac{1}{2}\) on both sides, so it comes down to showing the numerators are the same. \[6k^3+15k^2+11k+2\overbrace{=}^{\large ?}(k+1)(6(k+1)^2-3(k+1)-1)\]

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

Let's pull out a factor of \(k+1\) from the left side via long division: